OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin2x + sin x = 0

OpenStudy (anonymous):

can someone please help!:(

OpenStudy (amoodarya):

sin2x=- sinx sin 2x= sin(-x) 2x=-x +2k pi 2x=180-(-x) +2k pi

OpenStudy (amoodarya):

2x=-x+2k pi --->3x=2k pi x=2k pi /3 so x=0,2pi/3 ,4pi/3 and 2x=180+x+2k pi ----> x=180 +2k pi so just x=180 total roots in your interval are 0, 2pi/3, 4pi/3,pi

OpenStudy (anonymous):

@amoodarya what about pi/3?

OpenStudy (amoodarya):

sin(2)(pi/3) +sinpi/3 =sin 120 +sin 60 =sqrt3/2+sqrt3/2 =sqrt3 so pi/3 is not a root

OpenStudy (amoodarya):

sin 2x=2 sinx cosx so other solving 2 sinx cos x +sinx =0 sinx(2cos x+1)=0 so sinx =0 and cos x=-1/2

OpenStudy (anonymous):

a. x = 0, π, 4pi/3 b. x = 0, π,pi/3,2pi/3 c. x = 0, π, pi/3,5pi/3 d. x = 0, π, 3pi/2

OpenStudy (anonymous):

These are the option i have..

OpenStudy (amoodarya):